# A Framework for Double-Blind Federated Adaptation of Foundation Models

### v1

@@ Abstract
The availability of foundational models (FMs) pre-trained on large-scale data has advanced the state-of-the-art in many computer vision tasks. While FMs have demonstrated good zero-shot performance on many image classification tasks, there is often scope for performance improvement by adapting the FM to the downstream task. However, the data that is required for this adaptation typically exists in silos across multiple entities (data owners) and cannot be collated at a central location due to regulations and privacy concerns. At the same time, a learning service provider (LSP) who owns the FM cannot share the model with the data owners due to proprietary reasons. In some cases, the data owners may not have the resources to even store such large FMs. Hence, there is a need for algorithms to adapt the FM in a double-blind federated manner, i.e., the data owners do not know the FM or each other’s data and the LSP does not see the data for the downstream tasks. In this work, we propose a framework for double-blind federated adaptation of FMs using fully homomorphic encryption (FHE). The proposed framework first decomposes the FM into a sequence of FHE-friendly blocks through knowledge distillation. The resulting FHE-friendly model is adapted for the downstream task via low-rank parallel adapters that can be learned without backpropagation through the FM. Since the proposed framework requires the LSP to share intermediate representations with the data owners, we design a privacy-preserving permutation scheme to prevent the data owners from learning the FM through model extraction attacks. Finally, a secure aggregation protocol is employed for federated learning of the low-rank parallel adapters. Empirical results on four datasets demonstrate the practical feasibility of the proposed framework.
@@ 1 Introduction
Figure 1: Illustration of the double-blind federated adaptation.
@@
In the dynamic world of artificial intelligence (AI), the advent of foundational models (FMs) has marked a transformative era. Characterized by their vast scale, extensive pre-training, and versatility, FMs such as GPT [40, 4], CLIP [41], BERT [9, 15, 32], Stable Diffusion [43], Segment Anything [23], and Vision Transformers [10, 33] have advanced the state-of-the-art (SOTA) in many machine learning, computer vision, and language processing tasks. Though vision and multimodal FMs have demonstrated good zero-shot performance on many image classification tasks, there is often scope for performance improvement when faced with challenging out-of-domain tasks (e.g., medical images or satellite imagery). Hence, it becomes essential to adapt the FM for the downstream task.
@@
Adaptation of FMs for downstream tasks involves two main challenges - computational complexity and data availability. The simplest way to adapt an FM for a downstream classification task is via transfer learning - treat the FM as a feature extractor, append a classification head, and learn only the parameters of this classifier using the available training data (while the FM parameters remain frozen). If the classification head involves a single linear layer, it is referred to as linear probing. It is also possible to perform partial (only a selected subset of parameters are adapted) or full finetuning of the parameters of the FM based on the downstream data. Recently, two other broad categories of parameter-efficient fine-tuning (PEFT) approaches have been proposed for FM adaptation. The first approach is called prompt learning [20], where input (or intermediate) prompts are learned instead of modifying the FM parameters. The second approach is referred to as adapters [18, 8, 36], where additional components are added to the FM model architecture, and only these new parameters are learned. Adapters can be further categorized into sequential (e.g., low-rank adaptation a.k.a. LoRA [18]) and parallel (e.g., low-rank side adapter a.k.a. LoSA [36]) adapters. Except for transfer learning and parallel adapters, all the other adaptation techniques require partial or complete backpropagation of the gradients through the FM, which is computationally expensive. Finally, when only query access to a black-box FM is allowed, gradients can be estimated through zeroth-order optimization (ZOO). While ZOO avoids backpropagation, it is also computationally expensive because it requires numerous forward passes.
@@
The other major challenge in FM adaptation is that downstream training data required for adaptation may not be available to the learning service provider (LSP) who owns the FM. Moreover, this data may be distributed across multiple data owners (e.g., multiple hospitals or banks) and cannot be collated due to privacy concerns and regulations. This necessitates collaboration between the LSP and data owners to enable FM adaptation. Federated Learning (FL) [35] is a paradigm that promises to alleviate this data availability challenge by facilitating collaborative model training across distributed data repositories, while maintaining the privacy of the underlying data. FL encompasses a spectrum of applications, from healthcare [2, 51] and finance [34, 31] to video surveillance [48] and beyond [12]. It has two primary scenarios: cross-silo (few data owners) and cross-device FL (large number of data owners) [22], each catering to different scales of collaboration, thereby enabling both large and small-scale entities to benefit.
@@
In this work, we focus on the problem of federated adaptation of a foundation model (for an out-of-domain downstream image classification task) by an LSP (server) through collaboration with multiple data owners (clients) under two core constraints: (i) Model privacy - the LSP wants to retain full ownership of the FM and does not want to share the FM with the data owners; and (ii) Data privacy - the data owners do not want to leak their data to the LSP or other data owners. We jointly refer to these constraints as double-blind privacy. We make the following contributions:
@@
  • We propose a framework for double-blind federated adaptation of FMs based on well-known cryptographic constructs such as fully homomorphic encryption (FHE) and secure multi-party computation (MPC).
  • The proposed framework first distills the knowledge from the FM into a sequence of FHE-friendly transformer attention blocks. It then employs an interactive protocol based on a privacy-preserving permutation scheme to perform encrypted inference without leaking the model and the data. It also leverages a low-rank parallel adapter for local learning and an MPC protocol for FL aggregation.
  • We demonstrate the practical feasibility of the proposed framework through experiments on four datasets.
@@ 2 Related Work
Figure 2: Illustration of the block decomposition and non-linear functions that need to be approximated (highlighted in red).
@@
Foundation models: FMs have been highly successful in computer vision [10, 41, 33], natural language processing [40, 32, 52, 15, 7], and beyond [45, 42]. In particular, the two-stage training strategy has shown to be effective, where FMs are first pre-trained on a large dataset for general understanding, and then fine-tuned on a small downstream dataset to learn task-specific features. However, their vast scale introduces significant challenges, particularly in fine-tuning, that hinder their practical applicability.
@@
Private inference: The advent of machine learning as a service (MLaaS) has underscored the critical need for privacy-preserving techniques in ML, particularly in inference tasks. The concept of private inference (PI) has emerged as a pivotal solution to safeguard data and model privacy [37, 47, 46, 49, 57, 27, 19]. Vision transformers (ViTs), despite their superior performance in various tasks, pose significant challenges due to their computational complexity, especially when coupled with cryptographic constructs such as FHE [1, 11, 5] and MPC [13, 24] that are required for PI. The literature surrounding the optimization of PI frameworks primarily focuses on reducing computational and communication overheads while maintaining or improving model accuracy. SAL-ViT [57] is a framework designed to enhance the efficiency of PI on ViT models. Iron [14] introduced secure protocols aimed at optimizing matrix multiplication and several complex non-linear functions integral to transformer models, such as Softmax, GELU activations, and LayerNorm. An alternative approach is to develop PI-friendly transformer architectures and softmax approximations, such as MPCViT [53] and MPCFormer [27]. MPCViT introduces an MPC-friendly adaptation of ViTs that aims to balance the trade-off between accuracy and efficiency during inference in MPC settings. MPCFormer utilizes MPC and knowledge distillation to address latency and reduction in inference quality for transformers.
@@
Adaptation of foundation models: The primary issue in adapting FMs is their massive size, making it impractical for individual users or clients with limited computational resources to fine-tune or even store them. Various PEFT techniques such as adapters [17, 30], prompt learning [28], low-rank adaptation (LoRA) [18], and low-rank side adaptation [36] have been proposed. Numerous variants of LoRA, such as AdaLoRA [56], Delta-LoRA [60], IncreLoRA [54], and QLoRA [8] are also available. These methods specifically target the transformer attention blocks that form the core of these models. LoRA [18] introduces adjustments to the weight matrices within these blocks, thus enabling efficient fine-tuning with a reduced computational load. However, LoRA necessitates backpropagation through the backbone, thus increasing the total time taken to update the model. Examples of studies exploring PEFT techniques for LLMs with federated learning include [58] and [55].
@@ 3 Problem Formulation
Problem Statement: A task-specific classification head $\mathcal{H}_{\eta}$ and a side adapter $\mathcal{A}_{\theta}$ are appended to the FM to enable the adaptation. Thus, the adapted model ~$\widetilde{\mathcal{M}}_{\widetilde{\psi}}$ can be considered a combination of the original FM, the side adapter, and the classifier, i.e., ~$\widetilde{\mathcal{M}}_{\widetilde{\psi}}=(\mathcal{M}_{\psi}||\mathcal{A}_{\theta})\circ\mathcal{H}_{\eta}$, where ||||| | indicates that these functions operate in parallel and $\circ$ is the composition operator. The goal of the double-blind federated adaptation framework is to collaboratively learn the parameters $\eta$ and $\theta$ under the following constraints: (i) Data Privacy: the LSP does not learn anything about the local datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$ and data owner $\mathcal{P}_{k}$ does not learn anything about other local datasets $\mathcal{D}_{j}$, where $j\neq k$; (ii) Model Privacy: the data owners do not learn anything about the FM $\mathcal{M}_{\psi}$.
@@
Vanilla Federated Adaptation: Let $\mathcal{L}_{k}(\theta,\eta)=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbb{L}(\hat{y}_{i}^{k},y_{i}^{k})$ be the average loss at client $k$, where $\mathbb{L}$ denotes the per-sample loss and $\hat{y}_{i}^{k}=\widetilde{\mathcal{M}}_{\widetilde{\psi}}(\mathbf{x}_{i}^{k})$ is the prediction output by the adapted model. Federated adaptation can be posed as a distributed optimization problem [35], where the goal is to learn the global model parameters $(\theta,\eta)$ such that:
$\min_{(\theta,\eta)}~{}\sum_{k=1}^{K}\alpha_{k}\mathcal{L}_{k}(\theta,\eta),$
@@
where $\alpha_{k}=\frac{N_{k}}{\sum_{j=1}^{K}N_{j}}$. In each round $t$ of FL adaptation, the server broadcasts the previous model parameters $\left(\theta^{(t-1)},\eta^{(t-1)}\right)$. Each client computes the local model parameters $\left(\theta_{k}^{(t)},\eta_{k}^{(t)}\right)$ and these local updates are aggregated by the server to obtain the current global model parameters $\left(\theta^{(t)},\eta^{(t)}\right)$. For example, simple FedAvg aggregation function can be represented as follows:
$\left(\theta^{(t)},\eta^{(t)}\right)=\sum_{k=1}^{K}\alpha_{k}\left(\theta_{k}^{(t)},\eta_{k}^{(t)}\right).$ (1)
@@
Note that $t\in[1,T]$, where $T$ is the number of communication rounds and the model parameters $(\theta_{k}^{(0)},\eta_{k}^{(0)})$ for the first round are typically initialized randomly by the server.
@@
Challenges: The above vanilla federated adaptation is not privacy-preserving because it requires computation of $\hat{y}_{i}^{k}=\widetilde{\mathcal{M}}_{\widetilde{\psi}}(\mathbf{x}_{i}^{k})$, where the core FM $\mathcal{M}_{\psi}$ is available only at the server and the local training datasets are available only with the respective clients. Hence, it is essential to design a mechanism for computing $\mathcal{M}_{\psi}(\mathbf{x}_{i}^{k})$ without violating the data and model privacy constraints. Moreover, the sharing of local updates $\left(\theta_{k}^{(t)},\eta_{k}^{(t)}\right)$ with the server could also potentially leak information about the local datasets [59]. Hence, the aggregation step in Eq. 1 must be performed securely without revealing the local updates to the server.
4 Proposed Adaptation Framework
@@
The proposed framework for double-blind federated adaptation consists of the following steps: (1) FHE-friendly Distillation: Prior to federated adaptation, the server performs an offline distillation of the original FM into a FHE-friendly model using the auxiliary dataset. (2) Encrypted Inference: During adaptation, each client encrypts its local data using FHE and sends them to the server. The server and client engage in an interactive protocol to perform inference on the encrypted data. (3) Local Learning: Based on the intermediate outputs received from the server, each client locally updates the side network and classification head. (4) Secure Aggregation: The server securely aggregates these local updates through a multi-party computation (MPC) protocol with the clients. The overall training workflow of the proposed framework is summarized in Figure 2. At the end of this process, each client receives a copy of the global parameters of the side network and classifier. At the time of inference, the encrypted inference step (Step 2) is repeated. Based on the intermediate outputs received from the server, the client utilizes the global side adapter and classifier to obtain the final class prediction.
4.1 FHE-friendly Distillation
The first step in the proposed framework is to distill the knowledge available in the given FM into an FHE-friendly model. The server performs this step using the auxiliary dataset $\mathcal{D}_{aux}$. Assuming that the given FM follows a modular transformer encoder architecture as shown in Figure 3, we can consider the FM as a sequence of $L$ attention blocks. Let $\mathbf{b}_{\ell-1}$ be the input to the $\ell^{\text{th}}$ attention block $\mathcal{B}_{\psi_{\ell}}$ and $\mathbf{b}_{\ell}$ be the corresponding output. The objective of FHE-friendly distillation is to learn an FHE-friendly approximation of $\mathcal{B}_{\psi_{\ell}}$ denoted as ^$\widehat{\mathcal{B}}_{\widehat{\psi}_{\ell}}$ so that when the encrypted input $\mathcal{E}(\mathbf{b}_{\ell-1})$ is provided as input, the server can compute the encrypted output $\mathcal{E}(\mathbf{b}_{\ell})=\widehat{\mathcal{B}}_{\widehat{\psi}_{\ell}}(\mathcal{E}(\mathbf{b}_{\ell-1}))$. Here, $\mathcal{E}$ denotes the encryption operation.
@@ 4.1 FHE-friendly Distillation
Approximating Non-linear Functions: The primary challenge in encrypted inference is approximation of the non-linear functions because most well-known FHE schemes (e.g., CKKS [5]) allow only polynomial operations. A transformer attention block has three key non-linear operations: Softmax, GELU, and LayerNorm. In this work, Softmax is approximated using a Taylor series approximation of the exponential function $(e^{x})$:
@@
followed by normalization through division by the sum of the calculated exponential values. The error bound of this approximation is the remainder term, e.g. $\dfrac{e^{\xi}}{(d+1)!}x^{d+1}$, for some $\xi$ between 00 and $x$. Furthermore, GELU activation is approximated via a simple quadratic function:
@@
The LayerNorm function and Softmax require a division step, which is implemented based on the iterative protocol described in [6] following the Goldschmidt’s algorithm:
@@
Distilling the FM: After approximating all nonlinearities (Softmax, GELU, and Inverse) in the FM and developing an FHE-friendly FM, we enhance the performance of the FHE-friendly model through knowledge distillation. Initially, we apply layerwise distillation by aligning hidden representations at four positions: (i) the embedding layer, (ii) the attention matrix within each transformer block (considering raw values before normalization), (iii) hidden states after each transformer block, and (iv) the final prediction layer [21, 16, 27]. During distillation, we dedicate the first half of the training epochs to distilling the embedding and transformer blocks (both the attention matrix and hidden states) and then proceed to distill the prediction layer, following the approach in [21].
4.2 Encrypted Inference
Since FMs often have a very deep neural network architecture, it is challenging to perform encrypted inference over the full FM based on currently available FHE constructs that support only limited multiplicative depth. Attempting to do so will result in large approximation errors, which may be tolerable for simple inference tasks, but can potentially derail the federated learning process. One potential solution is to apply bootstrapping at regular intervals along the computational path. However, bootstrapping is computationally very expensive and makes the framework practically infeasible (especially under the thin client assumption). Hence, we propose performing encrypted inference only over a single transformer attention block at a time. The output of this block is sent back to the client, which decrypts the intermediate result, performs re-encryption of the intermediate result $\mathcal{E}(\mathcal{F}(\mathcal{E}(\mathbf{b}_{\ell})))$, and returns it to the server. Here, $\mathcal{F}$ denotes the decryption operation.
The downsides of performing encrypted inference over one attention block at a time are two-fold. Firstly, it increases the communication cost because encrypted intermediate outputs are exchanged between the client and the server after every block in every FL round. Since communication efficiency is not one of our core constraints, we consider the increased communication cost as a limitation, not a red flag. However, since the intermediate representations $\mathbf{b}_{\ell}$ after every attention block are accessible to the client in plaintext form, a malicious client could use $(\mathbf{b}_{\ell-1},\mathbf{b}_{\ell})$ pairs for multiple training samples to mount a model extraction attack [29] and learn the parameters of each transformer block. This clearly violates the model privacy constraint.
To circumvent the above problem and preserve model privacy, we leverage the fact that each communication round in the federated adaptation phase involves a batch of samples. Let $\mathcal{E}(\mathbf{B}_{\ell})=[\mathcal{E}(\mathbf{b}_{\ell}^{1}),\mathcal{E}(\mathbf{b}_{\ell}^{2}),\cdots,\mathcal{E}(\mathbf{b}_{\ell}^{n})]$ be a batch of encrypted intermediate representations corresponding to a client, where $n$ is the batch size. Before sending these representations to the client, the server applies a $n\times n$ permutation matrix $\mathbf{\Pi}_{\ell}$ and sends only the permuted batch $\mathcal{E}(\mathbf{B}_{\ell})\cdot\mathbf{\Pi}_{\ell}=[\mathcal{E}(\mathbf{b}_{\ell}^{\pi(1)}),\mathcal{E}(\mathbf{b}_{\ell}^{\pi(2)}),\cdots,\mathcal{E}(\mathbf{b}_{\ell}^{\pi(n)})]$ to the client. Here, $[\pi(1),\pi(2),\cdots,\pi(n)]$ represents a random permutation of $[1,2,\cdots,n]$. This permutation matrix $\mathbf{\Pi}_{\ell}$ can be randomly selected for each block $\ell$ in each communication round. Thus, the client never sees corresponding pairs $(\mathbf{b}_{\ell-1}^{i},\mathbf{b}_{\ell}^{i})$ for any training sample $i$ in the batch, ensuring protection against model extraction attacks.

### v2

@@ Abstract
Foundation models (FMs) excel in zero-shot tasks but benefit from task-specific adaptation. However, privacy concerns prevent data sharing among multiple data owners, and proprietary restrictions prevent the learning service provider (LSP) from sharing the FM. In this work, we propose BlindFed, a framework enabling collaborative FM adaptation while protecting both parties: data owners do not access the FM or each other’s data, and the LSP does not see sensitive task data. BlindFed relies on fully homomorphic encryption (FHE) and consists of three key innovations: (i) FHE-friendly architectural modifications via polynomial approximations and low-rank adapters, (ii) a two-stage split learning approach combining offline knowledge distillation and online encrypted inference for adapter training without backpropagation through the FM, and (iii) a privacy-boosting scheme using sample permutations and stochastic block sampling to mitigate model extraction attacks. Empirical results on four image classification datasets demonstrate the practical feasibility of the BlindFed framework, albeit at a high communication cost and large computational complexity for the LSP. The code can be found at https://github.com/tnurbek/blindfed.
@@ 1 Introduction
Foundation models (FMs) have transformed artificial intelligence, achieving state-of-the-art results across machine learning, computer vision, and natural language processing. Prominent examples include GPT [46, 7], CLIP [47], BERT [12, 18, 37], Stable Diffusion [49], Segment Anything [28], and Vision Transformers [13, 38]. Though vision and multimodal FMs have demonstrated good zero-shot performance, there is often scope for performance improvement when faced with challenging out-of-domain tasks (e.g., medical images or satellite imagery). Hence, it becomes essential to adapt the FM for the downstream task.
@@
Adaptation of FMs for downstream tasks involves two main challenges: computational complexity and data availability. The simplest approach is transfer learning, where the FM serves as a frozen feature extractor, and only a classification head is trained – linear probing is this head is a single linear layer. It is also possible to perform partial (only a selected subset of parameters are adapted) or full finetuning of the parameters of the FM based on the downstream data. Recent parameter-efficient fine-tuning (PEFT) methods fall into two categories: (i) prompt learning [23], which learns input or intermediate prompts without modifying FM parameters, and (ii) adapters [21, 11, 41], which add trainable components to the FM. Adapters include sequential (e.g., low-rank adaptation a.k.a. LoRA [21]) and parallel (e.g., low-rank side adaptation a.k.a. LoSA [41]) variants. Except for transfer learning and parallel adapters, all the other adaptation techniques require partial or complete backpropagation through the FM, which is computationally expensive. Zeroth-order optimization (ZOO) offers a backpropagation-free alternative for black-box FMs but incurs high cost due to numerous forward passes.
Figure 1: Conceptual illustration of BlindFed framework for double-blind federated adaptation of a foundation model.
@@
The other major challenge in FM adaptation is the unavailability of downstream training data to the learning service provider (LSP) who owns the FM. Moreover, this data may be distributed across multiple data owners (e.g., multiple hospitals or banks) and cannot be collated due to privacy concerns and regulations. Thus, FM adaptation requires collaboration between the LSP and data owners. Federated learning (FL) [40] addresses this challenge by enabling collaborative training across entities while preserving data confidentiality. FL has been applied in many applications [5, 59, 39, 36, 54, 15, 3] and operates mainly in two settings: cross-silo (few data owners) and cross-device FL (large number of data owners) [26].
@@
In this work, we focus on cross-silo federated adaptation of an FM (for an out-of-domain downstream image classification task) by an LSP (server) through collaboration with multiple data owners (clients) under two core constraints: (i) Model privacy - the LSP wants to retain full ownership of the FM and does not want to share the FM with the data owners; and (ii) Data privacy - clients do not want to reveal their data to the LSP or to each other. We jointly refer to these constraints as double-blind privacy (see Figure 1). We make the following four contributions:
@@
  • We propose the BlindFed framework for double-blind federated adaptation of FMs based on well-known cryptographic tools such as fully homomorphic encryption (FHE) and secure multiparty computation (MPC).
  • We modify the given FM into an FHE-friendly architecture, leveraging existing ideas such as polynomial approximations and low-rank parallel adapters.
  • We propose a two-stage split learning approach, where the FHE-friendly FM blocks are first pre-trained via offline knowledge distillation, followed by online encrypted inference to train local parallel adapters, with MPC-based secure aggregation for the global adapter.
  • For stronger model privacy, we introduce a privacy boosting scheme based on sample-level permutation and stochastic block sampling.
@@ 2 Related Work
Figure 2: Overview of the proposed BlindFed framework for double-blind federated adaptation. The framework consists of three main components: (1) FHE-friendly architecture redesign, where the original foundation model (FM) is modified by approximating non-linear operations; (2) offline distillation, where the approximated blocks are fine-tuned via knowledge distillation using an auxiliary dataset; and (3) online adaptation, where clients interact the FHE-enabled FM under homomorphic encryption, performing local updates on the parallel adapter and classification head.
@@
Foundation models: FMs have been highly successful in computer vision [13, 47, 38], natural language processing [46, 37, 61, 18, 10], and beyond [51, 48]. In particular, the two-stage training strategy has shown to be effective, where FMs are first pre-trained on a large dataset for general understanding and then fine-tuned on a small downstream dataset to learn task-specific features. However, their vast scale introduces significant challenges, particularly in fine-tuning, that hinder their practical applicability.
@@
Private inference: The advent of machine learning as a service (MLaaS) has underscored the need for privacy-preserving techniques in ML, particularly in inference tasks. The concept of private inference (PI) has emerged as a pivotal solution to safeguard data and model privacy [42, 53, 52, 55, 67, 32, 22]. While vision transformers (ViTs) achieve strong performance, their high computational cost makes PI challenging, especially under cryptographic techniques such as FHE [1, 14, 8] and MPC [16, 29]. Most PI literature focuses on reducing computational and communication overheads while preserving accuracy. SAL-ViT [67] improves ViT efficiency in PI, whereas Iron [17] optimizes matrix multiplication and key non-linear transformer operations (Softmax, GELU, LayerNorm). Another direction involves PI-friendly transformer designs, such as MPC-ViT [62], which adapts ViTs for MPC with an accuracy-efficiency trade-off, and MPCFormer [32], which combines MPC and knowledge distillation to reduce latency and maintain inference quality.
@@
Adaptation of foundation models: The primary issue in adapting FMs is their massive size, making it impractical for individual users or clients with limited computational resources to fine-tune or even store them. Various PEFT techniques such as adapters [20, 35], prompt learning [33], low-rank adaptation (LoRA) [21], and low-rank side adaptation (LoSA) [41] have been proposed. Numerous variants of LoRA, such as AdaLoRA [66], Delta-LoRA [70], IncreLoRA [63], QLoRA [11], LoRA-GA [58], and LoFT [56] further optimize adaptation efficiency. These methods specifically target the transformer attention blocks, with LoRA modifying weight matrices to enable efficient fine-tuning with a lower computational load. However, LoRA still requires backpropagation through the backbone, increasing the total time taken to update the model. PEFT in federated settings has also been explored for LLMs [68, 64].
@@ 3 Problem Formulation
Problem Statement: Let $\widetilde{\mathcal{M}}_{\widetilde{\psi}}$ denote the FM adapted for the downstream task. The goal of the BlindFed framework is to collaboratively learn the parameters $\widetilde{\psi}$ under the following constraints: (i) Data Privacy: the LSP does not learn anything about the local datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$ and the data owner $\mathcal{P}_{k}$ does not learn anything about other local datasets $\mathcal{D}_{j}$, where $j\neq k$; (ii) Model Privacy: the data owners do not learn anything about the original FM $\mathcal{M}_{\psi}$.
@@
Vanilla Federated Adaptation: Let $\mathcal{L}_{k}(\widetilde{\psi})=\frac{1}{N_{k}}\sum_{i=1}^{N_{k}}\mathbb{L}(\hat{y}_{i}^{k},y_{i}^{k})$ be the average loss at client $k$, where $\mathbb{L}$ denotes the per-sample loss and $\hat{y}_{i}^{k}=\widetilde{\mathcal{M}}_{\widetilde{\psi}}(\mathbf{x}_{i}^{k})$ is the prediction output by the adapted model. Federated adaptation can be posed as a distributed optimization problem [40], where the goal is to learn the global model parameters $\widetilde{\psi}$ such that:
[…]
@@
where $\alpha_{k}=\frac{N_{k}}{\sum_{j=1}^{K}N_{j}}$. In each round $t$ of FL adaptation, the server broadcasts the previous model parameters $\widetilde{\psi}^{(t-1)}$. Each client computes the local model parameters $\widetilde{\psi}_{k}^{(t)}$, and these local updates are aggregated by the server to obtain the current global model parameters $\widetilde{\psi}^{(t)}$. For example, simple FedAvg aggregation function can be represented as follows:
$\widetilde{\psi}^{(t)}=\sum_{k=1}^{K}\alpha_{k}\widetilde{\psi}_{k}^{(t)}.$ (2)
@@
Note that $t\in[1,T]$, where $T$ is the number of communication rounds and the model parameters $\widetilde{\psi}_{k}^{(0)}$ for the first round are typically initialized randomly by the server.
@@
Challenges: The above vanilla federated adaptation is not privacy-preserving because it requires computation of $\hat{y}_{i}^{k}=\widetilde{\mathcal{M}}_{\widetilde{\psi}}(\mathbf{x}_{i}^{k})$, where the core FM $\mathcal{M}_{\psi}$ is available only at the server and the local training datasets are available only with the respective clients. Hence, it is essential to design a mechanism for computing $\widetilde{\mathcal{M}}_{\widetilde{\psi}}(\mathbf{x}_{i}^{k})$ without violating the data and model privacy constraints. Moreover, the sharing of local updates $\widetilde{\psi}_{k}^{(t)}$ with the server could also potentially leak information about the local datasets [69]. Hence, the aggregation step in Eq. 2 must be performed securely without revealing the local updates to the server.
4 Proposed BlindFed Framework
@@
The BlindFed framework for double-blind federated adaptation (Figure 2) comprises two components: (1) FHE-friendly Architecture Redesign – The FM is first modified into an FHE-friendly model by approximating non-linear operations and adding a classification head as well as a parallel adapter; (2) Two-stage Split Learning – In the first offline stage, the approximated individual blocks are fine-tuned through knowledge distillation from the original FM using an auxiliary dataset. In the second online stage, clients encrypt their local data using FHE and interact with the server to perform encrypted inference. Based on intermediate outputs from the server, clients locally update the parallel adapter and classification head. The server then uses an MPC-based secure aggregation to combine these updates into global parameters, which are shared back with all clients. The overall training workflow of the proposed framework is summarized in Figure 3. During inference, the encrypted inference step is repeated. Based on the intermediate outputs received from the server, the client utilizes the global parallel adapter and classifier to obtain the final class prediction. To enhance model privacy, the server can further incorporate sample-level permutations and stochastic block sampling.
4.1 FHE-friendly Architecture Redesign
[…]
Figure 3: FHE-friendly architecture redesign. Each transformer block’s non-linear operations – GELU activations, Softmax attention, and the division step in LayerNorm – are replaced with low-degree polynomial approximations (denoted “Quad” for GELU and “ASoftmax” for Softmax). A lightweight parallel adapter and classification head are then trained on the client side.
@@ 4.1 FHE-friendly Architecture Redesign
Approximating Non-linear Functions: Encrypted inference is limited to polynomial operations in most FHE schemes (e.g., CKKS [8]), requiring polynomial approximations for non-linear functions in transformer blocks: Softmax, GELU, and LayerNorm. In this work, Softmax is approximated using a Taylor series approximation of the exponential function $(e^{x})$:
@@
followed by normalization through division by the sum of the calculated exponential values. The error bound of this approximation is the remainder term, e.g. $\frac{e^{\xi}}{(d+1)!}x^{d+1}$, for some $\xi$ between 0 and $x$. Furthermore, GELU activation is approximated via a simple quadratic function:
@@
The LayerNorm function and Softmax require a division, which is implemented via Goldschmidt’s algorithm [9]:
@@
[…]
4.2 Two-stage Split Learning
In the re-designed FHE-friendly FM, only the server stores the approximated FM; each client keeps a local parallel adapter and classifier. Training proceeds in two stages.
Stage 1: Offline Distillation. Before any collaboration, the server trains the approximated FM (student) from the original FM (teacher) on the auxiliary dataset $\mathcal{D}_{aux}$. After replacing all non-linearities (Softmax, GELU, and Inverse) with their approximations, we distill four types of representations: (i) embeddings, (ii) attention matrices (pre-normalization), (iii) hidden states after each block, and (iv) final prediction layer [24, 19, 32]. Following [24], the first half of epochs distills (i)-(iii); the second half distills (iv). Details of the distillation process appear in Appendix C.
Stage 2: Online Adaptation. This step is performed via an interactive protocol between the clients and the server, which can be further divided into three phases: (i) encrypted inference, (ii) local learning, and (iii) secure aggregation.
4.2.1 Encrypted Inference
FMs exceed the multiplicative depth supported by current FHE schemes; evaluating the whole network homomorphically would incur large approximation errors or require frequent (and impractical) bootstrapping, especially under the thin client assumption. Hence, we propose performing encrypted inference only over a single transformer block at a time. After each block, the client decrypts and re-encrypts the intermediate representation $\mathcal{E}(\mathcal{F}(\mathcal{E}(\mathbf{b}_{\ell})))$, and returns it back to the server. Here, $\mathcal{F}$ is the decryption operation.